# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ A.1 Proof of Theorem 3
For each $C_{\alpha}$, (8) reveals that the smallest data-dependent level $\widetilde{\alpha}$ choice for which $Z_{n+1}$ is excluded is $1/\mathcal{E}^{Z^{n}}(Z_{n+1})$. This implies,
$\frac{\mathbb{I}\{Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\leq\frac{\mathbb{I}\{Z_{n+1}\not\in C_{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}\}}{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}=\mathcal{E}^{Z^{n}}(Z_{n+1}),$
@@
$\mathrm{E}_{\widetilde{\alpha}}^{P}$ $\left[\frac{P(Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right$]
@@
$=\mathrm{E}^{P}\left[\frac{\mathbb{I}\{Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\right]\leq\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
@@ A.3 Proof of Proposition 2
By Proposition 3 in Koning (2025a), the conditional likelihood ratio is the restriction of
@@
to $O$. By definition of $\overline{Q}$, we have, for every event AA,

### v2

@@ A.1 Proof of Theorem 3
For each $\overline{C}_{\alpha}$, (8) reveals that the smallest data-dependent level $\widetilde{\alpha}$ choice for which $Z_{n+1}$ is excluded is $1/\mathcal{E}^{Z^{n}}(Z_{n+1})$. This implies,
$\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\leq\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}\}}{1/\mathcal{E}^{Z^{n}}(Z_{n+1})}=\mathcal{E}^{Z^{n}}(Z_{n+1}),$
@@
$\mathrm{E}_{\widetilde{\alpha}}^{P}$ $\left[\frac{P(Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right$]
@@
$=\mathrm{E}^{P}\left[\frac{\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\}}{\widetilde{\alpha}}\right]\leq\mathrm{E}^{P}\left[\mathcal{E}^{Z^{n}}(Z_{n+1})\right]\leq 1.$
@@ A.2 Proof of Proposition 1
To prove Proposition 1, we first prove the following lemma.
Lemma 3.
If $Z^{n}$ is exchangeable under $Q$ then
$Q^{Z^{n}|O,Z_{n+1}}=\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}),$ (18)
@@
Proof of Proposition 1.
We have
$\textnormal{LR}^{|O}(z^{n+1})=\frac{dQ^{|O}}{d\textnormal{Unif}(O)}(z^{n+1})$
$=\frac{d(Q^{Z_{n+1}|O}\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}{d(\textnormal{Unif}[Z^{n+1}]\otimes\textnormal{UnifWR}([Z^{n+1}]\setminus\{Z_{n+1}\}))}(z^{n+1})$
$=\frac{dQ^{Z_{n+1}|O}}{d\textnormal{Unif}[Z^{n+1}]}(z_{n+1})=(n+1)\times Q^{Z_{n+1}|O}(z_{n+1})$
$=:\textnormal{LR}^{Z_{n+1}|O}(z_{n+1}).$
@@ A.3 Proof of Proposition 2
By Proposition 3 in koning2023post, the conditional likelihood ratio is the restriction of
@@
to $O$. By definition of $\overline{Q}$, we have, for every event $A$,